(10-6y+e^(-3x))dx-2dy=0 


1
Expert's answer
2021-11-01T19:46:26-0400
"(10-6y+e^{-3x})dx-2dy=0"

"y'+3y=\dfrac{1}{2}e^{-3x}+5"

Find the integration factor


"\mu=e^{3x}"

Then


"y'e^{3x}+3ye^{3x}=\dfrac{1}{2}+5e^{3x}"

"d(e^{3x}y)=(\dfrac{1}{2}+5e^{3x})dx"

Integrate


"\int d(e^{3x}y)=\int(\dfrac{1}{2}+5e^{3x})dx"

"e^{3x}y=\dfrac{1}{2}x+\dfrac{5}{3}e^{3x}+C"

"y=\dfrac{1}{2}xe^{-3x}+\dfrac{5}{3}+Ce^{-3x}"